After eight years of dating, a female broke up with her lover on a rainy day. They chose to split up in the same area where everything about them began. The boy is crying and rushing north at a speed of 5 feet per second, while the girl is strolling east at a speed of 1 foot per second, wondering if she made the right decision. 5 seconds after they started traveling to a new life without each other, how quickly are they detaching from each other?


1
Expert's answer
2021-10-29T04:55:40-0400

Distance traveled by the boy is "5t" feet. Distance traveled by the girl is "1t" feet.

By the Pythagorean Theorem distance "s" between them is


"s(t)=\sqrt{(5t)^2+(1t)^2}=\sqrt{26}t(ft)"

How quickly are they detaching from each other


"v(t)=s'(t)=(\sqrt{26}t)'=\sqrt{26} (ft/sec)"

They are detaching from each other at a spped of "\sqrt{26}" from each other.